OpenStudy (anonymous):

Determine whether the ordered pair is a solution to the given linear inequality: x-7y≥ -3 (a) (4,1) , (b) (9,2), (c) (0,0) , (d) (-16,-2)

OpenStudy (anonymous):

I would probably guess sleek-feathered ones.

OpenStudy (anonymous):

Here is what I got

OpenStudy (anonymous):

(a) 4 - 7(1) ≥ -3 4 - 7 ≥ -3 -3 ≥ -3 True = Solution (b) 9 - 7(2) ≥ -3 9 - 14 ≥ -3 -5 ≥ -3 False = Not a solution. (c) 0 - 7(0) ≥ -3 0 - 0 ≥ -3 0 ≥ -3 True = Solution (d) -16 - 7(-2) ≥ -3 -16 + 14 ≥ -3 -2 ≥ -3 True = Solution